OpenStudy (anonymous):

A boat takes 3.0hr to travel 28km down a river, then 6.0hr to return. How fast is the river flowing?

OpenStudy (lastdaywork):

Let v = speed of boat with respect to river V = speed of river with respect to shore t = time in downstream journey T = time in upstream journey s = distance Then you'll get two equations with two unknowns - (v+V) = s/t (v-V) = s/T Solve for V

OpenStudy (mrnood):

Once again, this question requires the student to make an assumption. The assumption is that the boat moves at a constant speed through the water. This should have been stated in the question, but otherwise the above response is correct.